Six years ago, you purchased a callable bond with fifteen years until maturity. The bond has a $1,000 par value and pays interest semiannually. The bond has 9% coupon rate and a 6% yield to maturity. The bond offers three years of call protection and a 2% call premium.

a. How much did you pay for the bond at the time of purchase?

b. Today, the firm called the bond. What is the bond’s yield to call?

c. Did the firm make the right decision to call the bond? Why or why not?
Answer #1

1.
=PV(6%/2,2*15,-1000*9%/2,-1000)=1294.00662024205

2.
=RATE(6*2,9%*1000/2,-PV(6%/2,2*15,-1000*9%/2,-1000),1000*(1+2%))*2=3.7777535979978%

3.
Yes as yield to call is less than yield to maturity